OpenStudy (anonymous):

The 14 students in Mateo's computer programming class are using the computer lab. Some computers are shared by pairs of students, and 2 students work alone. Which equation could you use to determine how many computers are shared in the computer lab? let c be the number of shared computers. (see reply bar for multiple choice options)

OpenStudy (anonymous):

\[A. \frac{ c }{ 2 } - 2 = 14 \] \[B. 2c - 2 = 14\] C. \[2c + 2 = 14 \] \[D. \frac{ c }{ 2 } + 2 = 14 \]

OpenStudy (anonymous):

C

OpenStudy (raden):

C too

OpenStudy (anonymous):

there are twice as many people as shared computers: people = 2c